Selling buying a home has become very difficult due to the economical crisis which has severely affected the housing sector in recent years. Home sellers participating in a survey indicated that it typically takes between 6 and 14 months to sell a house. Suppose that the distribution of the time that it takes to sell a house can be approximated by a normal distributed with a mean of 10 months and a population standard deviation of 3 months.

(a) What percent of the homes take less than 9 months to sell?

(b) What percent of the homes take between 10 and 16 months to sell?

(c) How long does it take for top 10% of the houses to be sold?

(d) If 25 sold homes were randomly selected, what is the probability that the sample mean is less than 9 months? Hint: Use central limit theorem.

(e) Which one is more likely to happen: Randomly selecting a home that takes less than 9 months to sell (part-a) or Finding a random sample of 25 homes with a mean selling time of less than 9 months (part-d). Compare parts a and d, and enter a or d in the box below.
Answer #1

a)

for normal distribution z score =(X-μ)/σ
here mean=       μ= 10
std deviation   =σ= 3.0000
probability = P(X<9) = P(Z<-0.33)= 0.3707~37.07%

b)

probability = P(10<X<16) = P(0<Z<2)= 0.9772-0.5= 0.4772~47.72%

c)

for 90th percentile critical value of z= 1.28
therefore corresponding value=mean+z*std deviation= 13.84 months

d)

sample size       =n= 25
std error=σ=σ/√n= 0.6000
probability = P(X<9) = P(Z<-1.67)= 0.0475~ 4.75%

e)

option A is correct